Write the explicit formula for a_n given a1=-7 and a common difference of 1/2
Answer by stanbon(75887)

Write the explicit formula for a(n) given
a(1)=-7 and a common difference of 1/2
----
a(n) = a(1) + (n-1)d
---
a(n) = -7 + (n-1)(1/2)
---
a(n) = -7 + (n/2)-(1/2)
---
a(n) = (-15/2) + (n/2)
============================
Cheers,
Stan H.
============================